Circular No. 8868

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2007 R2 (GIBBS) A. R. Gibbs reports his discovery of a comet on four co-added 30-

s unfiltered CCD images obtained with the 0.68-m Schmidt telescope in the course of the Catalina Sky Survey (“discovery” observation tabulated below), the comet showing a compact 7′′ coma and a short, narrow tail 10′′ long in p.a. 250^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, other CCD astrometric observers have also commented on the object’s cometary appearance. J. W. Young (Table Mountain 0.61-m f/16 reflector) notes that his images from Sept. 11.36–11.39 UT show a slightly elongated 8′′-diameter coma with a fairly bright condensation, and a very straight tail is 40′′ long in p.a. 255^{o} is seen in several long exposures; images from Sept. 12.5 show a 8′′ coma slightly elongated toward the 30′′-long tail (p.a. 255^{o}–260^{o}). J. E. McGaha (Tucson, AZ, 0.62-m f/5.1 reflector) reports that fived stacked 60-s images from Sept. 11.44-11.45 show a 30′′ coma elongated in p.a. 245^{o} with a very narrow tail extending 90′′ at this same position angle. P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m f/6 reflector) writes that his stacked images from Sept. 12.1 show a diffuse coma of diameter 4′′ with a 10′′ tail in p.a. 260^{o}.

2007 UT  α_{2000}  δ_{2000}  Mag.
Sept.10.41401  3^{h}06^{m}39.^{s}7  +17^{o}40′47′′  18.2

The available astrometry, the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2007-R55.

T = 2007 Aug. 26.359 TT  ω = 352.^{o}796
e = 0.55779  Ω = 8.509 }2000.0
q = 1.45571 AU  i = 1.410
a = 3.29193 AU n^{o} = 0.165017 P = 6.0 years

V455 ANDROMEDAE = HS 2331+3905 N. N. Samus, Institute of Astronomy, Moscow, informs us that this cat-

aclysmic variable, which has undergone its first known major outburst this past week (cf. CBET 1053), has been assigned the permanent variable-star designation V455 And. E. Broens, Mol, Belgium, reports additional visual magnitude estimates on behalf of the Werkgroep Veranderlijke Sterren, Vereniging Voor Sterrenkunde: Sept. 6.837 UT, 8.7 (A. Diepvens, Balen, Belgium); 7.883, 9.2 (Broens); 9.872, 9.9 (H. Hautecler, Boutsersem, Belgium); 10.809, 10.3 (D. Vansteelant, Varsenare, Belgium).

2007 September 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green